# Resilience of Quantum Teleportation Fidelity for Bipartite Mixed States near Schwarzschild and Dilaton Black Holes

### v2

@@ I Introduction:
The organization of these paper is as follows. In section II, we discuss the vacuum structure for Dirac fields in a Schwarzschild and GHS Dilaton spacetime. Section III discusses genuine tripartite entanglement of SLOCC class of states viz. $GHZ$ state and $W$ state and also studies the teleportation fidelities of bipartite counterparts generated from these tripartite states after tracing out parties. This is followed by conclusion in section $IV$.
@@ II.2.2 Tripartite states subjected to black holes:
In the figure 1, a schematic diagram of our model shows tripartite states, where each qubits of the state are held by Alice ($A$), Bob ($B$) and Cliff ($C$). There may or may not be genuine tripartite entanglement among parties as in the cases of $GHZ$ and $W$ states. The rectangular box with dashed border represent the physical system of the tripartite state. The yellow circular region surrounding the black hole depicts the event horizon. The left hand side region of the black hole shows the physical system of the tripartite state before the parties Bob and Cliff are exposed to event horizon whereas the right hand side of the black hole represent the physical system of the same tripartite state after Bob and Cliff are exposed to event horizon of the black hole. Throughout the transition Alice remains in the flat region.
@@ II.2.3 GHZ state in Black Hole
We consider the $GHZ$ state as is defined in eq.(29) and we assume that initially the qubits are distributed among parties Alice ($A$), Bob ($B$) and Cliff ($C$). Also we consider that Alice stays in the flat (stationary) region while Bob and Cliff are near the event horizon of (i)Schwarzschild black hole and (ii) Dilaton Black Hole. Under this scenario, using eqs.(12) and (13), the wave function of $GHZ$ state is given as
@@
Figure 2: We plot the tangle of the $GHZ$ state against monochromatic frequency ($\omega$) and Hawking temperature ($T$). Fig.(A) shows the variation of tangle $\tau(\rho_{GHZ}^{wfabc})$ when Hawking temperature ($T$) is varied from 0 to $5$ and monochromatic frequency ($\omega$) is varied from 0 to $10000$ in the background of Schwarzschild black hole. Fig.(B) shows the variation of same $\tau(\rho_{GHZ}^{wfabc})$ when charge ($Q$) is varied from 0 to $20$ and monochromatic frequency ($\omega$) is varied from 0 to $50$ in the backdrop of Dilaton black hole.
@@ II.3.1 Protoype $W$ state in Black hole:
We consider the $W$ state as is defined in eq.(30) and we assume that initially the qubits are distributed among parties Alice ($A$), Bob ($B$) and Cliff ($C$). Alice stays in the flat (stationary) region while Bob and Cliff are near the event horizon of these two black holes. Under this scenario, using eqs.(12) and (13), the wave function of $W$ state is given by
@@
The concurrence of the state $\rho_{W}^{wfac}$ (or $\rho_{W}^{wfac}$) are plotted against the monochromatic frequency ($\omega$) and Hawking temperature ($T$) in the following figure.
@@
Below we plot the teleportation fidelity of the state $\rho_{W}^{wfac}$ (or $\rho_{W}^{wfac}$) against the parameters $\omega$ and $T$.
@@ III.0.1 Non-Protoype $W$ state in Black holes:
Figure 9: Variation of concurrence of bipartite mixed states $\rho_{W_{1}}^{wfac}$ derived from tripartite non-prototype $W_{1}$ state in the background of Schwarzschild black hole. Here $T$ is varied from $1$ to $5$ and $\omega$ is varied from 0 to $1$, where fig.(b) shows 2D plot of teleportation fidelity with respect to Hawking temperature ($T$) and fig.(b) shows the same with respect to monochromatic frequency ($\omega$).
@@
Figure 10: Variation of concurrence of bipartite mixed states $\rho_{W_{1}}^{wfac}$ derived from tripartite non-prototype $W_{1}$ state in the background of Dilaton black hole. Here $D$ is varied from $0.1$ to $10$ and $\omega$ is varied from 0 to $1$, where fig.(b) shows 2D plot of teleportation fidelity with respect to monochromatic frequency ($\omega$) and fig.(c) shows the same with respect to Dilaton parameter ($D$).
@@ IV Conclusion:
We have investigated the feasibility of quantum teleportation in the vicinity of black holes by analyzing bipartite mixed states derived from tripartite GHZ and W-class entangled states. In our setup, two observers (Bob and Cliff) approach the event horizon of either a Schwarzschild or a GHS Dilaton black hole, while a third (Alice) remains in flat spacetime. After tracing out one party, we evaluated the resulting bipartite state’s suitability as a teleportation channel using teleportation fidelity as the benchmark. Our results demonstrate that, despite the degradation of bipartite entanglement (quantified by concurrence) due to gravitational effects and Hawking radiation, teleportation fidelity remains above the classical threshold of $\frac{2}{3}$ for channels derived from W-class states. In contrast, GHZ-derived channels fail to support teleportation due to their lack of residual bipartite entanglement. Specifically, for Schwarzschild black holes, teleportation fidelity lies between 0.715–0.745, while for Dilaton black holes it ranges between 0.709–0.712 across the studied parameter regimes of Hawking temperature $T$, monochromatic frequency $\omega$, and Dilaton parameter $D$. This indicates that quantum teleportation remains viable near black holes provided the initial tripartite state retains useful bipartite entanglement after partial tracing. The resilience of teleportation fidelity, even as concurrence diminishes, suggests that quantum communication protocols can withstand certain relativistic gravitational effects. Our study thus establishes a foundation for further exploration of quantum information processing in curved spacetime. Future work could extend this analysis to other tripartite or multipartite states, alternative black hole models (e.g., rotating or charged black holes), and other quantum information protocols such as quantum key distribution and dense coding. Investigating the mathematical relationship between teleportation fidelity and quantum-gravitational parameters also remains an open direction.

### v3

@@ I Introduction:
Prior studies on quantum teleportation process and its fidelities in black hole spacetimes [10, 36, 37, 38, 39] focused on bipartite pure states or tripartite systems, overlooking derived bipartite mixed states from tracing out horizon affected parties. In this work, we extend this in three ways: (i) comparing teleportation fidelity in Schwarzschild and GHS Dilaton black holes with identical tripartite states, revealing how the Dilaton parameter($D$) introduces qualitatively different degradation patterns; (ii) explicitly distinguish between prototype $W$ states and non-prototype $W_{1}$ states, demonstrating that their different entanglement structures lead to distinct teleportation fidelity behaviors; and (iii) showing fidelity exceeds classical thresholds for $W$ class states despite concurrence loss, over wide parameters. This understanding offers practical insights for quantum communication networks that may operate in strong gravitational fields.
The organization of this paper is as follows. In section II, we discuss the vacuum structure for Dirac fields in a Schwarzschild and GHS Dilaton spacetime. Section III discusses genuine tripartite entanglement of SLOCC class of states viz. $GHZ$ state and $W$ state and also studies the teleportation fidelities of bipartite counterparts generated from these tripartite states after tracing out parties. This is followed by conclusion in section $IV$.
@@ II.2.2 Tripartite states subjected to black holes:
In the figure 1, a schematic diagram shows tripartite states held by Alice (A, flat region), Bob (B), and Cliff (C) near Schwarzschild/Dilaton event horizon($r_{h}$). “Near-horizon” for B/C means fixed $r\gtrsim r_{h}$ ($r_{h}=2M$ Schwarzschild, $r_{h}=2M-2D$ Dilaton), where local Dirac modes mix via Hawking effect (eqs.(6)–(9), (18)–(21)). Particles propagate on outgoing null geodesics $e^{-i\omega(t-r_{*})}$, but entanglement computed at snapshot before full infall ($\tau_{\text{cross}}\sim r_{h}\ln[(r-r_{h})/r_{h}]$ proper time). There may or may not be genuine tripartite entanglement among parties as in the cases of $GHZ$ and $W$ states. The rectangular box with dashed border represent the physical system of the tripartite state. The yellow circular region surrounding the black hole depicts the event horizon. The left hand side region of the black hole shows the physical system of the tripartite state before the parties Bob and Cliff are exposed to event horizon whereas the right hand side of the black hole represent the physical system of the same tripartite state after Bob and Cliff are exposed to event horizon of the black hole. Throughout the transition Alice remains in the flat region.
@@ II.2.3 GHZ state in Black Hole
We consider the $GHZ$ state as is defined in eq.(29) and we assume that initially the qubits are distributed among parties Alice ($A$), Bob ($B$) and Cliff ($C$). Also we consider that Alice stays in the flat (stationary) region while Bob and Cliff are near the event horizon of (i)Schwarzschild black hole and (ii) Dilaton Black Hole.
To obtain the wave function in the curved spacetime background, we must map the states of the observers near the event horizon (Bob and Cliff) from the Minkowski basis to the Kruskal basis. This mapping is derived from the Kruskal mode expansions given in eqs. (12) and (13), where we express the Kruskal vacuum and excited states as superpositions of the Schwarzschild outer and inner horizon states. The construction of the full wavefunction follows a systematic substitution procedure: (i) The flat Minkowski vacuum state $|0\rangle$ for a party near the horizon is replaced by the Kruskal vacuum: $|0\rangle\rightarrow|0\rangle_{K}=\mu|0\rangle_{out}|0\rangle_{in}+\nu|1\rangle_{out}|1\rangle_{in}$, and (ii) The flat Minkowski excited state $|1\rangle$ for a party near the horizon is replaced by the Kruskal single-particle state: $|1\rangle\rightarrow|1\rangle_{K}=|1\rangle_{out}|0\rangle_{in}$. Applying these substitutions term by term to the original $GHZ$ state of eq. (29) yields:
$|GHZ\rangle_{ABC}$ $=$ $\frac{1}{\sqrt{2}}\big(|0\rangle_{A}|0\rangle_{B}|0\rangle_{C}+|1\rangle_{A}|1\rangle_{B}|1\rangle_{C}\big)$
$\xrightarrow{\text{substitution}}$ $\frac{1}{\sqrt{2}}\Big[|0\rangle_{A}\big(\mu|0\rangle_{B_{out}}|0\rangle_{B_{in}}+\nu|1\rangle_{B_{out}}|1\rangle_{B_{in}}\big)\otimes\big(\mu|0\rangle_{C_{out}}|0\rangle_{C_{in}}+\nu|1\rangle_{C_{out}}|1\rangle_{C_{in}}\big)$
$\quad+|1\rangle_{A}\big(|1\rangle_{B_{out}}|0\rangle_{B_{in}}\big)\otimes\big(|1\rangle_{C_{out}}|0\rangle_{C_{in}}\big)\Big].$
[…]
@@
Figure 2: We plot the tangle of the $GHZ$ state against monochromatic frequency ($\omega$) and Hawking temperature ($T$). Fig.(A) shows the variation of tangle $\tau(\rho_{GHZ}^{wfabc})$ when Hawking temperature ($T$) is varied from 0 to $5$ and monochromatic frequency ($\omega$) is varied from 0 to $10000$ in the background of Schwarzschild black hole. Fig.(B) shows the variation of same $\tau(\rho_{GHZ}^{wfabc})$ when charge ($Q$) is varied from 0 to $20$ and monochromatic frequency ($\omega$) is varied from 0 to $50$ in the background of Dilaton black hole.
@@ II.3.1 Protoype $W$ state in Black hole:
We consider the $W$ state as is defined in eq.(30) and we assume that initially the qubits are distributed among parties Alice ($A$), Bob ($B$) and Cliff ($C$). Alice stays in the flat (stationary) region while Bob and Cliff are near the event horizon of these two black holes. Following the same substitution procedure detailed in Section II.2.3 for the $GHZ$ state, we apply the Kruskal mode expansion to the prototype $W$ state of eq.(30). Replacing $|0\rangle_{B}$, $|1\rangle_{B}$, $|0\rangle_{C}$, and $|1\rangle_{C}$ with their Kruskal counterparts as given in eqs.(12) and (13), and expanding term by term, we obtain the wave function of $W$ state as:
@@
The concurrence of the state $\rho_{W}^{wfac}$ (or $\rho_{W}^{wfab}$) are plotted against the monochromatic frequency ($\omega$) and Hawking temperature ($T$) in the following figure.
@@
Below we plot the teleportation fidelity of the state $\rho_{W}^{wfac}$ (or $\rho_{W}^{wfab}$) against the parameters $\omega$ and $T$.
@@ III.0.1 Non-Protoype $W$ state in Black holes:
Figure 9: Variation of teleportation fidelity of bipartite mixed states $\rho_{W_{1}}^{wfac}$ derived from tripartite non-prototype $W_{1}$ state in the background of Schwarzschild black hole. Here $T$ is varied from $1$ to $5$ and $\omega$ is varied from 0 to $1$, where fig.(b) shows 2D plot of teleportation fidelity with respect to Hawking temperature ($T$) and fig.(b) shows the same with respect to monochromatic frequency ($\omega$).
@@
Figure 10: Variation of teleportation fidelity of bipartite mixed states $\rho_{W_{1}}^{wfac}$ derived from tripartite non-prototype $W_{1}$ state in the background of Dilaton black hole. Here $D$ is varied from $0.1$ to $10$ and $\omega$ is varied from 0 to $1$, where fig.(b) shows 2D plot of teleportation fidelity with respect to monochromatic frequency ($\omega$) and fig.(c) shows the same with respect to Dilaton parameter ($D$).
@@
Based on our analysis presented in figs. 3,4,5,6 and figs.7,8,9,10, we compare the behaviors of concurrences and fidelities of prototype $W$ state and non-prototype $W_{1}$ state. The non-prototype $W_{1}$ state shows higher maximum concurrence ($C\approx 0.46$ in Schwarzschild, $C\approx 0.70$ in Dilaton) compared to the prototype $W$ state ($C\approx 0.26$ in Schwarzschild, $C\approx 0.12$ in Dilaton), indicating stronger residual bipartite entanglement. One can also observe the prototype $W$ state maintains finite concurrence for $D<1.16$, while the $W_{1}$ state extends to $D<1.5$, showing greater resilience to the Dilaton parameter. The prototype $W$ state exhibits teleportation fidelity in the range $0.715<f^{T}<0.745$ for Schwarzschild and $0.709<f^{T}<0.712$ for Dilaton, with systematic variation against $T$ and $\omega$. The $W_{1}$ state maintains $f^{T}>0.7$ across both backgrounds when entanglement persists, but shows sharper fidelity degradation near the critical Dilaton value. The prototype $W$ state offers more reliable performance across wider parameter ranges, making it suitable for scenarios with uncertain black hole parameters. The $W_{1}$ state provides higher potential fidelity but requires more controlled conditions, particularly regarding the Dilaton parameter. These differences arise from the distinct entanglement distributions in the two $W$ class states, demonstrating that the internal structure of tripartite entanglement significantly influences the teleportation capability of derived bipartite channels under gravitational effects.
@@ IV Conclusion:
We have investigated the feasibility of quantum teleportation in the vicinity of black holes by analyzing bipartite mixed states derived from tripartite GHZ and W-class entangled states. In our setup, two observers (Bob and Cliff) approach the event horizon of either a Schwarzschild or a GHS Dilaton black hole, while a third (Alice) remains in flat spacetime. After tracing out one party, we evaluated the resulting bipartite state’s suitability as a teleportation channel using teleportation fidelity as the benchmark. Our results demonstrate that, despite the degradation of bipartite entanglement (quantified by concurrence) due to gravitational effects and Hawking radiation, teleportation fidelity remains above the classical threshold of $\frac{2}{3}$ for channels derived from W-class states. In contrast, GHZ-derived channels fail to support teleportation due to their lack of residual bipartite entanglement. Specifically, for Schwarzschild black holes, teleportation fidelity lies between 0.715–0.745, while for Dilaton black holes it ranges between 0.709–0.712 across the studied parameter regimes of Hawking temperature $T$, monochromatic frequency $\omega$, and Dilaton parameter $D$. This indicates that quantum teleportation remains viable near black holes provided the initial tripartite state retains useful bipartite entanglement after partial tracing. The resilience of teleportation fidelity, even as concurrence diminishes, suggests that quantum communication protocols can withstand certain relativistic gravitational effects. While this work focused on the scenario where two parties (B and C) are near the horizon, the framework can be readily adapted to other configurations. For instance, if only Bob is near the horizon while Alice and Cliff remain in flat spacetime, the degradation would be isolated to a single party. In that case, the bipartite channel between Alice and Cliff would be expected to retain higher fidelity than the channels studied here, offering a more robust pathway for teleportation. Comparing such asymmetric configurations could provide deeper insights into how the location of gravitational degradation affects the distribution of entanglement and the viability of quantum networks in relativistic settings. Our static analysis assumes stationary observers at fixed $r\gtrsim r_{h}$, capturing snapshot entanglement/fidelity without dynamical infall or LOCC signal delays. Realistically, continuous horizon approach degrades states during finite classical communication, potentially dropping $f^{T}<2/3$ mid protocol. Extending our analysis to dynamical geometries with Hawking backreaction and quantum fluctuations is a key future direction. This would assess if teleportation fidelity stays above classical thresholds during evaporation and how the finite teleportation window scales with initial parameters, illuminating quantum information-gravity interplay in realistic astrophysical scenarios. Our study thus establishes a foundation for further exploration of quantum information processing in curved spacetime. Future work could extend this analysis to other tripartite or multipartite states, alternative black hole models (e.g., rotating or charged black holes), and other quantum information protocols such as quantum key distribution and dense coding. Investigating the mathematical relationship between teleportation fidelity and quantum-gravitational parameters also remains an open direction.